# OMoE: Diversifying Mixture of Low-Rank Adaptation by Orthogonal Finetuning

### v1

@@ 4.2 Main results
Single-Task Setup. In this setup, we compare OMoE with the performance of LoRA, DoRA and MixLoRA by employing these methods for fine-tuning a single task. The comparison with other baselines is attached in the Appendix D. The experimental results on diverse datasets are presented in Table 1. We provide the number of tunable parameters in the third column, which shows that OMoE significantly reduces the number of experts and tunable parameters by approximately 75%. Table 1 also reveals that OMoE outperforms the baselines in most tasks, with other tasks achieve respectable performance. Additionally, in most datasets, the performance gap between OMoE-LoRA and OMoE-DoRA is smaller than that between LoRA and DoRA. This indicates that OMoE is not sensitive to underlying fine-tuning methods and effectively utilizes the model’s capacity.
Multi-Task Setup. Table  2 presents the results of LoRA, DoRA, MOELORA, MiLoRA, MixLoRA and OMoE with LLaMA2-7B in multi-task learning. In contrast to the single-task setup in Table  1, during multi-task learning, we mix training data from ARC, BoolQ, OBQA, and PIQA to train the model, followed by separate evaluations to investigate the generalization ability of each method. The results indicate that, compared to single-task learning, LoRA and DoRA exhibit degradation in average accuracy in multi-task learning. In this configuration, MOELORA and MiLoRA exhibit comparable performance, showing slight degradation. MixLoRA benefits from top-2 routing with 8 experts, outperforming the single-task setup, albeit at the cost of requiring more parameters. In contrast, OMoE achieves stable performance improvements under minimal parameter conditions.
@@
4.3 Ablation studies and further analysis
Comparisons under different budgets of tunable parameters. We vary the budget of tunable parameters for OMoE-LoRA by adjusting the values of LoRA rank $r=16$ to $\{2,4,8,32\}$. We also modify the tunable parameters of MiLoRA and MoELoRA as baselines, consistent with  Zhang et al. (2024). The experimental results on the BoolQ and PIQA tasks are presented in Fig. 3, while performance on other benchmarks is presented in the Appendix E. The results indicate that under varying budgets of tunable parameters, OMoE-LoRA consistently outperforms the baselines and demonstrates resource efficiency by significantly reducing tunable parameters. As $r$ increases from 2 to 4, 8, and 16, the model’s accuracy benefits from the continuous increase in parameters. When $r=32$, the performance of OMoE declines due to the difficulties in convergence under limited training steps.
Effect of k and routing strategies. In Table  1 and  2, we set the number of LoRA experts, k, to 2. To investigate the effect of $k$, we vary $k$ to {3, 4, 5, 6, 7} and explore the performance of OMoE under different routing strategies. For comparison, we use Soft Routing and Top-2 Routing as baselines to validate the performance improvements brought by OMoE. The average performance results for the multi-task setting([arc-e, arc-c, boolq, obqa]) are presented in Fig. 4. The results indicate that: (a) Our OMoE method consistently outperforms both LoRA and vanilla MoE routing strategies. (b) As shown in Fig. 4(a), OMoE with Soft Routing exhibits a declining trend as the number of experts increases, likely due to excessive orthogonality among experts leading to meaningless representations and reduction that hinder overall performance. OMoE achieves its best performance with two experts, significantly surpassing vanilla MoE. (c) As illustrated in Fig. 4(b), our OMoE also benefits slightly from Top-2 routing, demonstrating that expert representation diversity enhances model performance. Furthermore, OMoE substantially reduces tunable parameters while maintaining expert diversity, allowing it to effectively integrate with various routing strategies.
Visulization of expert representations. As shown in Fig. 4(a), in our investigation of the impact of $k$ on OMoE, we found that OMoE with soft routing suffers from larger $k$. This is contrary to our intuition, as greater diversity among experts should lead to more robust performance. We hypothesize that the presence of multiple orthogonal experts (for $k\geq 3$ introduces redundancy. To validate this hypothesis, we conducted fine-tuning experiments on the OBQA dataset, gradually increasing the number of experts in OMoE under soft routing and visualizing the representations of randomly selected expert layers in large models. As illustrated in Fig.  5, when $k=2$, the experts within OMoE benefit from orthogonal constraints, leading to diversified representations. However, the increase in the number of experts ($k\geq 3$) inevitably introduces redundancy and may generate meaningless representations, thereby limiting overall performance. As depicted in Fig. 4(a), using two orthogonal experts yields the best results, maximizing the model’s capacity and minimizing the computational complexity. We also visualized representations across different layers in LLMs, please refer to the Appendix F for more details.
@@ 4.3 Ablation studies and further analysis
Figure 5: The representations under different numbers of experts in OMoE are visualized using a uniform t-SNE transformation for 2D representation reduction. The representations are obtained from layer 16 of LLaMA-2 7B. The visualization of more layers is provided in Appendix G.
Analysis of Model Efficiency. To demonstrate the inference efficiency of our OMoE method, we compared the inference latency and GPU memory utilization of LoRA, DoRA, MixLoRA, and OMoE under two typical routing strategies on a batch of input instructions. We assess efficiency based on two metrics: (a) the inference time required for generating responses (ms). (b) the GPU memory cost. The results are presented in Table 3.
As illustrated in Table 3, OMoE (Soft Routing) significantly reduces memory cost and inference latency compared to the well-performing MixLoRA. Specifically, OMoE (Soft Routing) is 30.1% faster than MixLoRA while decreasing GPU memory cost by 698 MiB. Although OMoE (Top-2 Routing) slightly increases computational overhead, it sightly enhances models’ performance. The advantages of OMoE (Soft Routing) and the analysis can be summarized as follows:
  • OMoE (Soft Routing) reduces the number of LoRA experts to modify Transformer layers at each decoding step, making generating new tokens more efficient. Compared to top-2 routing, OMoE merely calls two experts, so even with the additional computational steps for orthogonality, the inference efficiency is still improved.
@@
Method Latency (ms) Extra Memory (MiB)
LoRA 2,096 +1,630
@@
Table 3: The inference latency and memory cost of the LLama-2 7B model for generating a batch of responses using different PEFT methods.
Analysis of Layer-wise Orthogonality. After demonstrating that orthogonal experts possess a superior capability for handling fine-grained information, a natural question arises: ”Which layers in LLMs require orthogonal experts the most?” To conduct a more detailed analysis of the role of orthogonal experts across different layers in LLMs, we take LLama-2 7b as an example and simplify the model into three levels: low (from layer 1 to 10), medium (from layer 11 to 20), and high (from layer 21 to 32), injecting OMoE into these levels. Specifically, we propose four types of layer-wise orthogonal expert configurations:
  • OMoE-$\bigtriangleup$ injects OMoE into low levels while medium and high level are vanilla MoE.
  • OMoE-$\bigtriangledown$ injects OMoE into high levels while low and medium level are vanilla MoE.
  • OMoE-$\diamond$ injects OMoE into medium levels while low and high level are vanilla MoE.
@@ 5 Conclusion and Discussion
This paper investigates the expert collapse in MoE. We analyze the benefits of diversity in MoE and develop OMoE, a resource-efficient MoE variant designed to promote expert diversity, which facilitates stable and efficient performance improvements. In OMoE, the Gram-Schmidt process is leveraged to ensure that the experts’ representations lie within the Stiefel manifold. This process enables OMoE to reduce representation redundancy when fine-tuning MoE. Additionally, OMoE keeps the learning objective unchanged by imposing constraints. Our qualitative analysis and experimental results highlight the benefits of OMoE in efficient fine-tuning for LLMs. Based on the performance of OMoE on supervised fine-tuning, we believe higher potential of OMoE can be exploited with Reinforcement Learning from Human Feedback (RLHF) in future study.
@@
Appendix A Related Works
Parameter-Efficient Fine-tuning: To enhance the performance of LLMs for downstream tasks, fine-tuning is an unavoidable step. However, direct full-parameter fine-tuning is time-consuming and memory-intensive, posing challenges for practical applications. Consequently, researchers are motivated to explore parameter-efficient fine-tuning (PEFT) techniques aimed at reducing the time and memory costs during fine-tuning. The core idea of Parameter-Efficient Fine-Tuning (PEFT) is to fine-tune only a small subset of model parameters while keeping the majority of pre-trained parameters unchanged. Representative methods include Adapters, Prefix-tuning, and Prompt-tuning. Adapters Zhang et al. [2023b]; He et al. [2021]; Hu et al. [2022], one of the earliest approaches, improve the adaptability of large models by optimizing their architecture through automatic search. BitFit Zaken et al. [2021] and $(\text{IA})^{3}$ Liu et al. [2022] link learnable vectors to the hidden states of various modules within the Transformer layers using either multiplication or addition, thereby facilitating efficient fine-tuning. Prefix-tuning Li and Liang [2021] adds embeddings to the hidden states of the Transformer model, while Prompt-tuning Liu et al. [2021]; Zhu et al. [2024] only adds embeddings to the input layer. LoRA Hu et al. [2021], which utilizes low-rank matrix decomposition of linear layer weights, has become one of the popular PEFT methods, with many works dedicated to improving LoRA. AdaLoRA Zhang et al. [2023a] investigates the parameter allocation of LoRA modules. DoRA Liu et al. [2024] decomposes pre-trained weights into magnitude and direction components to reduce the number of trainable parameters. lora+ Hayou et al. [2024] sets different but fixed learning rates for the adapter matrices A and B in LoRA to correct the suboptimalities of LoRA. rsLoRA Kalajdzievski [2023] modifies LoRA by using appropriate scaling factors to achieve a computation/performance trade-off during fine-tuning. Vera Kopiczko et al. [2023] randomly frozes initialized LoRA matrices and only finetunes a set of scaling vectors.
MoE variants: MoE variants utilize a combination of parallel LoRA modules, referred to as ”experts,” and a gating mechanism that selects which experts to activate based on the input data. This architecture achieves a better trade-off between performance and training cost, making it a popular choice for efficiently fine-tuning large language models. Mixlora Li et al. [2024] and Moelora Liu et al. [2023b] integrate Low-Rank Adaptation (LoRA) with MoE for multi-task learning, constructing a resource-efficient sparse model. Llava-mole Chen et al. [2024] adopts a sparse combination of LoRA experts to address data conflicts when handling instruction data from different domains. SCMoE leverages contrastive learning within an MoE framework to enhance parameter efficiency during fine-tuning processes for LLMs. HydraLoRA Tian et al. [2024] introduces an asymmetric structure to the LoRA framework without requiring domain expertise, improving both training and inference performance and efficiency. MiLoRA Zhang et al. [2024] employs a prompt-aware routing mechanism, allowing it to reuse these results for subsequent tokens, thereby reducing latency. Teamlora Lin et al. [2024] emphasizes collaboration and competition among experts, further enhancing the effectiveness of low-rank adaptation. Some works apply MoE architectures to the inference stage of large models to achieve better generalization. LoraHub Huang et al. [2023] and LLMAdapters Hu et al. [2023] adopt the MoE structure by training multiple adapters and strategically selecting adapter combinations based on the domain during inference, thereby improving the model’s performance in multi-task environments.
Appendix B Datasets
The detailed information about the datasets used in the experiments are presented in Table 6. All datasets are download from Hugging Face.
@@ Appendix B Datasets
Appendix C Experimental settings
@@
Table 9: The Overall comparison of different PEFT methods for single-task learning. Most of the results are extracted from the original papers or reproduced by running the provided source code. The backbone model is LlaMA-2 7B.
@@
Table 10: Accuracy comparison of OMoE-LoRA with varying ranks for LLaMA2-7B on various benchmarks.
Method Rank $r$ Params (M) ARC-e ARC-c BoolQ PIQA Average
OMoE-LoRA 2 8.2 73.4 50.5 72.8 82.4 69.8
4 16.1 75.3 51.6 73.7 83.7 71.1
8 31.9 77.2 52.1 73.1 83.8 71.6
16 63.4 79.3 56.6 73.5 84.5 73.5
32 126.6 79.7 57.7 72.5 82.8 73.2
Computing infrastures We run all our experiments on NVIDIA A6000 (48GB) GPUs and setup with Python 3.10 and Ubuntu 20.04 on x86-64 CPUs.
@@
Hyper-parameters for OMoE In our experiments, unless otherwise specified, we set hyperparameter as illustrated in Table LABEL:Hyperparameter_OMoE. Under the above setting, OMoE will introduce 0.73% tunable parameters to the LLaMA-2 7B backbone.
Hyper-parameters for LoRA/DoRA and MixLoRA/MixDoRA In our experiments, we set hyperparameter for LoRA/DoRA and MixLoRA/MixDoRA as illustrated in Table 8, consistent with the settings in Li et al. [2024].
Hyper-parameters for other baselines In our experiments, we set hyperparameter for other baselines as follows, consistent with the settings in Zhang et al. [2024]:
  • P-tuning V2: The number of prompt tokens at each layer is set to 16, and the soft prompts are initialized with dimension 640, and then is projected to dimension 4096.
@@
Appendix D Additional Results on other baselines
In the main paper, we compared OMoE with three widely recognized and well-performing baselines: LoRA, DoRA, and MixLoRA, across the LLama 7B, 8B, and 13B models. Here, in addition to the three baselines presented in Table  1, we provide results from experiments involving 11 additional strong baselines on the Llama 7B backbone in Table  9. The results indicate that OMoE demonstrates significant improvements in both parameter efficiency and performance compared to the baselines.
Appendix E The robustness of OMoE under varying ranks and different numbers of tunable parameters
We vary the budget of tunable parameters for OMoE by modifying the values of rank r = 2, 4, 8, 16, 32. We also vary the MOELoRA method’s tunable parameter numbers. Futhermore, we can verify the robustness of OMoE under varing ranks, which is illustrated in Table 10. To achieve a better trade-off between performance and training costs, we set the hyperparameter $r$ to 16.
Appendix F Additional visualization of representations using different methods

### v2

@@ 4.2 Main Results
Single-Task Setup. In this setup, we compare OMoE with the performance of LoRA, DoRA and MixLoRA by employing these methods for fine-tuning a single task. The comparison with other baselines is provided in the Appendix C. The experimental results on diverse datasets are presented in Table 1. We provide the number of tunable parameters in the third column, which shows that OMoE significantly reduces the number of experts and tunable parameters by approximately 75%. Table 1 also reveals that OMoE outperforms the baselines in most tasks, with other tasks achieving respectable performance. Additionally, in most datasets, the performance gap between OMoE-LoRA and OMoE-DoRA is smaller than that between LoRA and DoRA. This indicates that OMoE is not sensitive to underlying fine-tuning methods and effectively utilizes the model’s capacity.
Multi-Task Setup. Table 2 presents the results of LoRA, DoRA, MoELoRA, MiLoRA, MixLoRA and OMoE with LLaMA-2 7B in multi-task learning. In contrast to the single-task setup in Table  1, during multi-task learning, we mix training data from ARC, BoolQ, OBQA, and PIQA to train the model, followed by separate evaluations to investigate the generalization ability of each method. The results indicate that, compared to single-task learning, LoRA and DoRA exhibit degradation in average accuracy in multi-task learning. In this configuration, MoELoRA and MiLoRA exhibit comparable performance, showing slight degradation. MixLoRA benefits from top-2 routing with 8 experts, outperforming the single-task setup, albeit at the cost of requiring more parameters. In contrast, OMoE achieves stable performance improvements under minimal parameter conditions. We also add experiments under a multi-task setting on LLaMA-3 8B, with the relevant results shown in the Table 3.
4.3 Ablation Studies and Further Analysis
Comparisons under Different Budgets of Tunable Parameters. We vary the budget of tunable parameters for OMoE-LoRA by adjusting the values of LoRA rank $r=16$ to $\{2,4,8,32\}$. We also modify the tunable parameters of MiLoRA and MoELoRA as baselines, consistent with  [43]. The experimental results on the BoolQ and PIQA tasks are presented in Fig. 3. Furthermore, we can verify the robustness of OMoE under varying ranks, which is illustrated in Table 4. To achieve a better trade-off between performance and training costs, we set the hyperparameter $r$ to 16. The results indicate that under varying budgets of tunable parameters, OMoE-LoRA consistently outperforms the baselines and demonstrates resource efficiency by significantly reducing tunable parameters. As $r$ increases from 2 to 4, 8, and 16, the model’s accuracy benefits from the continuous increase in parameters. When $r=32$, the performance of OMoE declines due to the difficulty in convergence under limited training steps.
Effect of k and Routing Strategies. In Tables 1 and  2, we set the number of LoRA experts, k, to 2. To investigate the effect of $k$, we vary $k$ to {3, 4, 5, 6, 7} and explore the performance of OMoE under different routing strategies. For comparison, we use Soft Routing and Top-2 Routing as baselines to validate the performance improvements brought by OMoE. The average performance results for the multi-task setting ([ARC-e, ARC-c, BoolQ, OBQA]) are presented in Fig. 4. The results indicate that:
(a) Our OMoE method consistently outperforms both LoRA and vanilla MoE routing strategies.
(b) As shown in Fig. 4(a), OMoE with Soft Routing exhibits a declining trend as the number of experts increases. The underlying reason is that the weighted averaging mechanism of soft routing proves incompatible with specialized and diversified expert modules. The performance degradation of OMoE in soft routing precisely corroborates OMoE’s effectiveness in promoting expert diversity. The robustness of soft routing when varying the number of experts demonstrates the existence of representational redundancy in vanilla MoE.
(c) As illustrated in Fig. 4(b), our OMoE also benefits slightly from Top-2 routing, demonstrating that expert representation diversity enhances model performance. Furthermore, OMoE substantially reduces tunable parameters while maintaining expert diversity, allowing it to effectively integrate with various routing strategies.
@@ 4.3 Ablation Studies and Further Analysis
Visulization of Expert Representations. OMoE can be regarded as a direct solution for achieving expert diversity in MoE architectures, accomplished by enforcing representation orthogonality. To visualize expert diversity, we conducted fine-tuning experiments on the OBQA dataset: gradually increasing the number of experts in OMoE under top-k routing while visualizing representations from randomly selected expert layers in large models. As illustrated in Fig. 5, the visualization results demonstrate OMoE’s successful achievement of expert diversity. Furthermore, as shown in Fig. 4(a), using two orthogonal experts achieves the optimal balance simultaneously achieving expert diversity while maintaining a minimal number of experts. We also visualized representations across different layers in LLMs and please refer to the Appendix D for more details.
@@
Figure 5: The representations under different numbers of experts in OMoE using top-2 routing are visualized via t-SNE. The representations are obtained from layer 16 of LLaMA-2 7B. The visualization of more layers is provided in Appendix E.
Analysis of Model Efficiency. To demonstrate the inference efficiency of our OMoE method, we compared the inference latency, GPU memory utilization and training time of LoRA, DoRA, MixLoRA, and OMoE under two typical routing strategies on a batch of input instructions. We assess efficiency based on three metrics: (a) the inference time required for generating responses (ms). (b) the GPU memory cost. (c) the training time for multi-task setting. The results are presented in Table 5.
As illustrated in Table 5, OMoE (Soft Routing) significantly reduces memory cost and inference latency compared to the well-performing MixLoRA. Specifically, OMoE (Soft Routing) is 30.1% faster than MixLoRA while decreasing GPU memory cost by 698 MiB. In terms of training time, OMoE requires 2.1 hours on the RTX A6000 in the multi-task setting (arc-e, arc-c, boolq, obqa), while MixLoRA takes 2.2 hours. Although OMoE (Top-2 Routing) slightly increases computational overhead, it slightly enhances models’ performance. The advantages of OMoE (Soft Routing) and the analysis can be summarized as follows:
  • OMoE (Soft Routing) reduces the number of LoRA experts to modify Transformer layers at each decoding step, making generating new tokens more efficient. Compared to OMoE (Top-2 Routing), OMoE merely calls two experts, so even with the additional computational steps for orthogonality, the inference efficiency is still improved.
@@
  • The orthogonalization in OMoE does slightly increase the computational overhead, but compared to MoE methods like MixLoRA, OMoE does not require training a router and does not need to calculate the router balancing loss, which reduces the computational burden. Overall, OMoE reduces the training time.
Method Latency (ms) Memory (MiB) Training time (h)
LoRA 2,096 +1,630 1.8h
@@
Table 5: The inference latency and memory cost of the LLaMA-2 7B model for generating a batch of responses using different baselines.
Analysis of Layer-Wise Orthogonality. After demonstrating that orthogonal experts possess a superior capability for handling fine-grained information, a natural question arises: "Which layers in LLMs require orthogonal experts the most?" To conduct a more detailed analysis of the role of orthogonal experts across different layers in LLMs, we take LLaMA-2 7B as an example and simplify the model into three levels: low (from layer 1 to 10), medium (from layer 11 to 20), and high (from layer 21 to 32), injecting OMoE into these levels. Specifically, we propose four types of layer-wise orthogonal expert configurations:
  • OMoE-$\bigtriangleup$ injects OMoE into low levels while medium and high levels are vanilla MoE.
  • OMoE-$\bigtriangledown$ injects OMoE into high levels while low and medium levels are vanilla MoE.
  • OMoE-$\diamond$ injects OMoE into medium levels while low and high levels are vanilla MoE.
@@
4.4 Discussion: Redundancy in Visualization
As illustrated in Fig. 5, when $k=2$, the experts within OMoE benefit from orthogonal constraints, leading to diversified representations. However, the increase in the number of experts ($k\geq 3$) inevitably introduces redundancy. To explain this phenomenon, we propose the following interpretation: The similar representations are the result of the final optimization objective, that is, in the learning process, the objective function being optimized forces the latter orthogonal experts to output representations similar to those of the previous orthogonal experts. For example: Firstly, we have two learned expert representations $(v_{1}=(1,0,0)$ and $(v_{2}=(0,0,5)$. Then, the newly added orthogonal expert $(v_{3}=(0,0.9,0)$ attempts to optimize along dim=2, but the final optimization objective forces it to approach the meaningful representation vector $(v_{1}=(1,0,0)$. Therefore, although the experts are orthogonal, expert 3 is still similar to expert 1 in the visualization, resulting in redundancy. This redundancy is caused by the optimization objective, which suppresses the diversity of more experts, rather than the inherent redundancy of the MoE.
5 Related Works
Parameter-Efficient Fine-Tuning: To enhance the performance of LLMs for downstream tasks, fine-tuning is an unavoidable step. However, direct full-parameter fine-tuning is time-consuming and memory-intensive, posing challenges for practical applications. Consequently, researchers are motivated to explore parameter-efficient fine-tuning (PEFT) techniques aimed at reducing the time and memory costs during fine-tuning. The core idea of Parameter-Efficient Fine-Tuning (PEFT) is to fine-tune only a small subset of model parameters while keeping the majority of pre-trained parameters unchanged. Representative methods include Adapters, Prefix-tuning, and Prompt-tuning. Adapters [45, 10, 14], one of the earliest approaches, improve the adaptability of large models by optimizing their architecture through automatic search. BitFit [41] and $(\text{IA})^{3}$ [27] link learnable vectors to the hidden states of various modules within the Transformer layers using either multiplication or addition, thereby facilitating efficient fine-tuning. Prefix-tuning [24] adds embeddings to the hidden states of the Transformer model, while Prompt-tuning [30, 46] only adds embeddings to the input layer. LoRA [13], which utilizes low-rank matrix decomposition of linear layer weights, has become one of the popular PEFT methods, with many works dedicated to improving LoRA. AdaLoRA [44] investigates the parameter allocation of LoRA modules. DoRA [29] decomposes pre-trained weights into magnitude and direction components to reduce the number of trainable parameters. Lora+ [9] sets different but fixed learning rates for the adapter matrices A and B in LoRA to correct the suboptimalities of LoRA. rsLoRA [18] modifies LoRA by using appropriate scaling factors to achieve a computation/performance trade-off during fine-tuning. Vera [20] randomly freezes initialized LoRA matrices and only fine-tunes a set of scaling vectors.
MoE Variants: MoE variants utilize a combination of parallel LoRA modules, referred to as "experts", and a gating mechanism that selects which experts to activate based on the input data. This architecture achieves a better trade-off between performance and training cost, making it a popular choice for efficiently fine-tuning large language models. Mixlora [22] and Moelora [28] integrate Low-Rank Adaptation (LoRA) with MoE for multi-task learning, constructing a resource-efficient sparse model. Loramoe [8] incorporates a localized balancing constraint, which enables two expert groups to maintain specialized functions between groups while achieving balanced workloads within each group. Llava-mole [4] adopts a sparse combination of LoRA experts to address data conflicts when handling instruction data from different domains. SCMoE [37] leverages contrastive learning within an MoE framework to enhance parameter efficiency during fine-tuning processes for LLMs. HydraLoRA [39] introduces an asymmetric structure to the LoRA framework without requiring domain expertise, improving both training and inference performance and efficiency. MiLoRA [43] employs a prompt-aware routing mechanism, allowing it to reuse these results for subsequent tokens, thereby reducing latency. Teamlora [25] emphasizes collaboration and competition among experts, further enhancing the effectiveness of low-rank adaptation. Some works apply MoE architectures to the inference stage of large models to achieve better generalization. LoraHub [16] and LLMAdapters [15] adopt the MoE structure by training multiple adapters and strategically selecting adapter combinations based on the domain during inference, thereby improving the model’s performance in multi-task environments.
@@ 6 Conclusion and Discussion
This paper investigates the expert collapse in MoE. We analyze the benefits of diversity in MoE and develop OMoE, a resource-efficient MoE variant designed to promote expert diversity, which facilitates stable and efficient performance improvements. In OMoE, the Gram-Schmidt process is leveraged to ensure that the experts’ representations lie within the Stiefel manifold. This process enables OMoE to reduce representation redundancy when fine-tuning MoE. Additionally, OMoE keeps the learning objective unchanged by imposing constraints. Our qualitative analysis and experimental results highlight the benefits of OMoE in efficient fine-tuning for LLMs. Based on the performance of OMoE on supervised fine-tuning, we believe the higher potential of OMoE can be exploited with Reinforcement Learning from Human Feedback (RLHF) in future study.
{ack}
This work was supported in part by the National Natural Science Foundation of China under Grant 62322316, the Key Research Project of Chinese Academy of Sciences under Grant No. KGFZD-145-25-21-01, the Open Fund of National Key Laboratory of Information Systems Engineering (No. 6142101240203), and the Beijing Nova Program under Grant 20220484077 and 20230484435.
@@
Appendix A Datasets
The detailed information about the datasets used in the experiments is presented in Table 8. All datasets are downloaded from Hugging Face.
@@ Appendix A Datasets
Appendix B Experimental settings
@@
Table 11: An Overall comparison of different PEFT methods for single-task learning is provided. Most of the results are extracted from the original papers or reproduced by running the provided source code. The backbone model is LLaMA-2 7B.
@@
Computing infrastructure We run all our experiments on NVIDIA A6000 (48GB) GPUs and set up with Python 3.10 and Ubuntu 20.04 on x86-64 CPUs.
@@
Hyper-parameters for OMoE In our experiments, unless otherwise specified, we set hyperparameters as illustrated in Table LABEL:Hyperparameter_OMoE. Under the above setting, OMoE will introduce 0.73% tunable parameters to the LLaMA-2 7B backbone.
Hyper-parameters for LoRA/DoRA and MixLoRA/MixDoRA In our experiments, we set hyperparameters for LoRA/DoRA and MixLoRA/MixDoRA as illustrated in Table 10, consistent with the settings in [22].
Hyper-parameters for other baselines In our experiments, we set hyperparameter for other baselines as follows, consistent with the settings in [43]:
  • P-tuning V2: The number of prompt tokens at each layer is set to 16, and the soft prompts are initialized with dimension 640, and then are projected to dimension 4096.
@@
Appendix C Additional Results on Other Baselines
In the main paper, we compared OMoE with three widely recognized and well-performing baselines: LoRA, DoRA, and MixLoRA, across the LLaMA-2 7B, 8B, and 13B models. Here, in addition to the three baselines presented in Table 1, we provide results from experiments involving 11 additional strong baselines on the Llama 7B backbone in Table 11. The results indicate that OMoE demonstrates significant improvements in both parameter efficiency and performance compared to the baselines.
Appendix D Additional Visualization of Representations using Different Methods